# Regular BTZ black holes from an infinite tower of corrections

### v1

@@ I Introduction
In this paper, we build upon the work done in Refs. [53, 54] (see also Refs. [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] for related work), which first explored the theories emerging from the dimensional regularization of the Gauss-Bonnet term [46, 47, 48, 49, 50, 51] to $2+1$ dimensions. Notably, these works found that the resulting black hole solutions closely resemble those of the higher-dimensional Einstein-Gauss-Bonnet theory [68, 69, 70], and represent Bañados-Teitelboim-Zanelli (BTZ) black holes [71, 72, 73] corrected by the higher-derivative terms. Following Ref. [44], we extend the dimensional regularization procedure to all Lovelock orders in $2+1$ spacetime dimensions. In the limit where an infinite tower of regularized Lovelock corrections is included, we obtain everywhere regular BTZ black holes. Remarkably, these black holes feature an extremal inner horizon, which prevents the classical mass inflation instabilities that typically afflict most regular black hole models [74, 75, 76]. Thus, beyond ensuring a fully regular geometry, the higher-derivative corrections also naturally evade mass inflation instabilities. This stands in stark contrast to most other $2+1$-dimensional black holes, where modifications of GR generally introduce singularities or non-extremal Cauchy horizons. [53, 54, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96].
@@ II Scalar-Tensor Theories from Regularized Lovelock Gravity
$\begin{array}[]{|c|c|c|c|c|}\hline\cr c_{n}&G_{2}-2/L^{2}&G_{3}&G_{4}-1&f(r)\\ \hline\cr\hline\cr 1&-\frac{8\ell^{2}X^{2}}{(1-2\ell^{2}X)^{2}}&8\ell^{2}X\frac{1-\ell^{2}X}{(1-2\ell^{2}X)^{2}}&\frac{\ell^{2}X}{-1+2\ell^{2}X}-\frac{3}{2}\sqrt{2\ell^{2}X}\tanh^{-1}\left(\sqrt{2\ell^{2}X}\right)&\frac{r^{2}f_{\rm BTZ}}{r^{2}-\ell^{2}f_{\rm BTZ}}\\ \hline\cr\frac{1}{n}&\frac{4X}{-1+2\ell^{2}X}-\frac{2}{\ell^{2}}\log(1-2\ell^{2}X)&\frac{4\ell^{2}X}{1-2\ell^{2}X}&-\sqrt{2\ell^{2}X}\tanh^{-1}\left(\sqrt{2\ell^{2}X}\right)&\frac{r^{2}}{\ell^{2}}\left(\exp\left(\frac{\ell^{2}}{r^{2}}f_{\rm BTZ}\right)-1\right)\\ \hline\cr\frac{1-(-1)^{n}}{2n}&\frac{4X}{-1+4\ell^{4}X^{2}}+\frac{2}{\ell^{2}}\tanh^{-1}(2\ell^{2}X)&\frac{8\ell^{4}X^{2}}{1-4\ell^{4}X^{2}}&\sqrt{\frac{\ell^{2}X}{2}}\left(\tan^{-1}\left(\sqrt{2\ell^{2}X}\right)-\tanh^{-1}\left(\sqrt{2\ell^{2}X}\right)\right)&\frac{r^{2}}{\ell^{2}}\tanh\left(\frac{\ell^{2}}{r^{2}}f_{\rm BTZ}\right)\\ \hline\cr\frac{\Gamma(\frac{n}{2})\delta_{0,k}}{\sqrt{\pi}\Gamma(\frac{n+1}{2})}&-\frac{16\ell^{4}X^{3}}{(1-4\ell^{4}X^{2})^{3/2}}&\frac{2}{(1-4\ell^{4}X^{2})^{3/2}}&{}_{2}F_{1}\left(-\frac{1}{4},\frac{3}{2},\frac{3}{4},4\ell^{4}X^{2}\right)&\frac{r^{2}f_{\rm BTZ}}{\sqrt{r^{4}+\ell^{4}f_{\rm BTZ}^{2}}}\\ \hline\cr\end{array}$
@@
$\mathcal{L}^{(n)}=\lim_{d\to 2n}\frac{\sqrt{-g}\mathcal{R}^{(n)}-\sqrt{-\tilde{g}}\tilde{\mathcal{R}}^{(n)}}{d-2n},$ (2)
@@
where $\{c_{n}\}$ are dimensionless coupling constants, $\ell$ is a new length scale, and we consider a negative cosmological constant $\Lambda=-1/L^{2}$, defined in terms of the anti de Sitter (AdS) radius $L$. When the series in Eq. (5) is truncated at $n=2$, the resulting theory coincides, up to field and coupling redefinitions, with the one studied in Refs. [53, 54]. The action (5) belongs to the class of $3D$ Horndeski theories (see Refs. [103, 104] for the full field equations) and possesses a shift symmetry in the scalar field, meaning it remains invariant under $\phi\to\phi+C$ for any constant $C$. As a result, there exists a conserved Noether current $j^{\mu}$, whose divergence-free condition, $\nabla_{\mu}j^{\mu}=0$, is equivalent to the scalar field equation. The explicit form of the Noether current can be determined using the results of Ref. [105].
@@ III Regular BTZ Black Holes and a Birkhoff Theorem
where $\mu$ is an integration constant. For the solution to describe a black hole, $\mu$ must be positive, and it is associated with the black hole’s mass. When $\mu>0$, the spacetime exhibits a conical-like singularity in its causal structure at $r=0$, resembling the behavior found in the Taub-NUT spacetime [73].
@@
where $f_{\rm BTZ}(r)$ is defined in Eq. (7), and $c_{1}=1$. The BTZ solution is trivially recovered when the series is truncated at $n=1$, and the solution of Refs. [53, 54] when truncated at $n=2$. Notoriously, Eq. (10) is exactly the equation that determines planar black holes in higher-dimensional Lovelock gravity [106, 107].
@@
However, this does not imply uniqueness. In particular, the BTZ geometry remains a solution to the field equations when the scalar field is constant. This mirrors the behavior of purely metric modified gravity theories in $2+1$ dimensions, which always admit the BTZ black hole as a solution [89]. This further reinforces the interpretation of Eq. (5) as a higher-curvature theory.
@@
where $f_{\rm BTZ}$ is given in Eq. (7). The metric is regular everywhere and free of singularities as long as $L>\ell$. Near the origin, it behaves as $f(r)=-r^{2}/\ell^{2}+\mathcal{O}(r^{4})$. In Fig. 1, we plot this metric function for a specific set of parameters, showing that the metric remains regular throughout. Other metric functions listed in Tab. 1 exhibit qualitatively similar shapes and regularity properties. It is important to note that all solutions are symmetric under $r\to-r$. The event horizon in all cases is located at $r_{+}=L\sqrt{\mu}$, similar to the BTZ black hole. Asymptotically, the spacetime of the regular black hole behaves like AdS, with $f(r)\sim-r^{2}/L_{\rm eff}^{2}$, where the effective AdS radius is given by $L_{\rm eff}^{2}=L^{2}-\ell^{2}$ for the metric in Eq. (15).
@@ IV Geodesics and tidal forces
$-\epsilon=-f(r)\dot{t}^{2}+\frac{\dot{r}^{2}}{f(r)}+r^{2}\dot{\varphi}^{2},$ (19)
@@
Geodesics do not terminate at $r=0$; instead, they cross the inner-extremal horizon and move toward negative values of $r$ [109], engaging in a cyclic motion between antipodal points in causally disconnected universes. This behavior is illustrated in Fig. 2 (bottom), where we show one complete cycle of a timelike geodesic.
@@ V Rotating Black Holes
The determinant of the induced metric on constant-$r$ hypersurfaces is given by $-r^{2}f(r)$. As a result, the horizon structure remains unchanged, with the event horizon located at $r_{+}=L\sqrt{\mu}$ and the inner-extremal horizon at $r_{-}=0$ in this coordinate system. These black holes also possess ergoregions when $g_{tt}>0$, where the Killing vector $\partial_{t}$ becomes spacelike.
@@ VI Conclusions
The regular black holes obtained in this work feature an inner-extremal horizon with zero surface gravity. An important avenue for future research would be to investigate their dynamical stability against mass inflation. Additionally, these black holes satisfy a Birkhoff-type theorem, implying that the exterior region of any matter configuration where the metric is of the form given in Eq. (6) will be described by the corresponding static black hole solution. This raises the interesting question of whether these regular black holes can form dynamically. Exploring gravitational collapse scenarios similar to those studied in Refs. [35, 36] could provide valuable insights into this possibility.

### v2

@@ I Introduction
In this paper, we build upon the work done in Refs. [53, 54] (see also Refs. [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] for related work), which first explored the theories emerging from the dimensional regularization of the Gauss-Bonnet term [46, 47, 48, 49, 50, 51] to $2+1$ dimensions. Notably, these works found that the resulting black hole solutions closely resemble those of the higher-dimensional Einstein-Gauss-Bonnet theory [68, 69, 70], and represent Bañados-Teitelboim-Zanelli (BTZ) black holes [71, 72, 73] corrected by the higher-derivative terms. Following Ref. [44], we extend the dimensional regularization procedure to all Lovelock orders in $2+1$ spacetime dimensions. In the limit where an infinite tower of regularized Lovelock corrections is included, we obtain everywhere regular BTZ black holes. Remarkably, these black holes feature an extremal inner horizon, which prevents the classical mass inflation instabilities that typically afflict most regular black hole models [74, 75, 76]. Thus, beyond ensuring a fully regular geometry, without singularities of any kind, the higher-derivative corrections also naturally evade mass inflation instabilities. This stands in stark contrast to the cases of i) the static BTZ black hole of GR, which is geodesically incomplete [77], and the rotating BTZ black hole which has a standard Cauchy horizon, being prone to mass inflation instability [78]; and ii) most other $2+1$-dimensional black holes, where modifications of GR generally introduce curvature singularities or non-extremal Cauchy horizons [53, 54, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98].
@@ II Scalar-Tensor Theories from Regularized Lovelock Gravity
[…]
@@
$\sqrt{-g}\mathcal{L}^{(n)}=\lim_{d\to 2n}\frac{\sqrt{-g}\mathcal{R}^{(n)}-\sqrt{-\tilde{g}}\tilde{\mathcal{R}}^{(n)}}{d-2n},$ (2)
@@
where we consider a negative cosmological constant $\Lambda=-1/L^{2}$, defined in terms of the anti de Sitter (AdS) radius $L$, $\{c_{n}\}$ are dimensionless coupling constants, and $\ell$ is a new length scale where beyond-GR effects become important, and that limits the spacetime curvature for the regular black hole solutions discussed in Sec. III. When the series in Eq. (5) is truncated at $n=2$, the resulting theory coincides, up to field and coupling redefinitions, with the one studied in Refs. [53, 54]. The action (5) belongs to the class of $3D$ Horndeski theories (see Refs. [105, 106] for the full field equations) and possesses a shift symmetry in the scalar field, meaning it remains invariant under $\phi\to\phi+C$ for any constant $C$. As a result, there exists a conserved Noether current $j^{\mu}$, whose divergence-free condition, $\nabla_{\mu}j^{\mu}=0$, is equivalent to the scalar field equation. The explicit form of the Noether current can be determined using the results of Ref. [107].
@@ III Regular BTZ Black Holes and a Birkhoff Theorem
where $\mu$ is an integration constant. For the solution to describe a black hole, $\mu$ must be positive, and it is associated with the black hole’s mass. When $\mu>0$, the spacetime exhibits a conical-like singularity in its causal structure at $r=0$, resembling the behavior found in the Taub-NUT spacetime [73]. The static BTZ black hole is thus geodesically incomplete [77].
@@
where $f_{\rm BTZ}(r)$ is defined in Eq. (7), and $c_{1}=1$. The BTZ solution is trivially recovered when the series is truncated at $n=1$, and the solution of Refs. [53, 54] when truncated at $n=2$. Notoriously, Eq. (10) is exactly the equation that determines planar black holes in higher-dimensional Lovelock gravity [108, 109]. This correspondence arises because the class of theories considered in Eq. (5), obtained via the limiting procedure in Eq. (2), can also be derived through a regularized Kaluza-Klein reduction of Lovelock gravity with a flat internal space [49, 50]. In this dimensional reduction, the scalar field $\phi$ corresponds to the logarithm of the scale factor of the internal space. Both the $2+1$-dimensional black holes discussed here and higher-dimensional planar black holes possess a planar (i.e., flat) base manifold geometry, with a scale factor given by $r$. By adopting the scalar field profile in Eq. (8) and considering a planar base manifold, we preserve these symmetries of the original higher-dimensional Lovelock theory prior to the reduction, and we expect the same functional form for the solutions.
@@
However, this does not imply uniqueness. In particular, the BTZ geometry remains a solution to the field equations when the scalar field is constant. This mirrors the behavior of purely metric modified gravity theories in $2+1$ dimensions, which always admit the BTZ black hole as a solution [91].
@@
where $f_{\rm BTZ}$ is given in Eq. (7). The metric is regular everywhere and free of singularities of any kind, as long as $L>\ell$. Near the origin, it behaves as $f(r)=-r^{2}/\ell^{2}+\mathcal{O}(r^{4})$. In Fig. 1, we plot this metric function for a specific set of parameters, showing that the metric remains regular throughout. Other metric functions listed in Tab. 1 exhibit qualitatively similar shapes and regularity properties. It is important to note that all solutions are symmetric under $r\to-r$. The event horizon in all cases is located at $r_{+}=L\sqrt{\mu}$, similar to the BTZ black hole. Asymptotically, the spacetime of the regular black hole behaves like AdS, with $f(r)\sim-r^{2}/L_{\rm eff}^{2}$, where the effective AdS radius is given by $L_{\rm eff}^{2}=L^{2}-\ell^{2}$ for the metric in Eq. (15).
@@
We emphasize that, despite some qualitative similarities between the solutions presented here and those in Refs. [22, 23], the underlying theories, spacetime geometries, equations of motion, and mechanism curing the singularity are fundamentally different. In particular, the theory defined by Eq. (5) ensures second-order field equations, in contrast to the higher-order equations generally present in Refs. [22, 23].
@@ IV Geodesics and tidal forces
$\epsilon=-f(r)\dot{t}^{2}+\frac{\dot{r}^{2}}{f(r)}+r^{2}\dot{\varphi}^{2},$ (19)
@@
Geodesics do not terminate at $r=0$, contrarily to the standard BTZ black hole [77]; instead, they cross the inner-extremal horizon and move toward negative values of $r$ [111], engaging in a cyclic motion between antipodal points in causally disconnected universes. This behavior is illustrated in Fig. 2 (bottom), where we show one complete cycle of a timelike geodesic.
@@ V Rotating Black Holes
The determinant of the induced metric on constant-$r$ hypersurfaces is given by $r^{2}f(r)$. As a result, the horizon structure remains unchanged, with the event horizon located at $r_{+}=L\sqrt{\mu}$ and the inner-extremal horizon at $r_{-}=0$ in this coordinate system. These black holes also possess ergoregions when $g_{tt}>0$, where the Killing vector $\partial_{t}$ becomes spacelike.
We note that the inner horizon of these regular solutions is extremal only due to the presence of the length scale $\ell$, as the rotating BTZ black hole is recovered in the limit $\ell\to 0$. Thus, in this case, the same mechanism that resolves the singularity also prevents mass inflation instabilities, even though these typically arise at much lower curvature scales than the singularity.
@@ VI Conclusions
The regular black holes obtained in this work feature an inner-extremal horizon with zero surface gravity. An important avenue for future research would be to investigate their dynamical stability against mass inflation; we expect no instability exists. Additionally, these black holes satisfy a Birkhoff-type theorem, implying that the exterior region of any matter configuration where the metric is of the form given in Eq. (6) will be described by the corresponding static black hole solution. This raises the interesting question of whether these regular black holes can form dynamically. Exploring gravitational collapse scenarios similar to those studied in Refs. [35, 36] could provide valuable insights into this possibility.